Two cyclists,Elvis and James start at the same time and in the same place.Elvis travels due east at 9 km per hour and James due north at 12 km per hour.How far are they from each other after one hour?
Answer by checkley77(12844)

A^2+B^2=C^2
9^2+12^2=C^2
81+144=C^2
C^2=228
C=SQRT228
C=15.1 MILES IS THE DISTANCE APART AFTER 1 HOUR.